Six types of workers are employed in each of the two workshops but at different rates of wages as follows Type of worker Workshop A Rate of wages per worker Rupees Mechanic Fitter Electrician Carpenter Smith Clerk 2 250 350 400 300 300 200 14 20 Workshop B No of workers No of workers Rate of wages per worker Rupees 300 18 300 50 425 7 350 12 10 6 350 1 500 2 In which of the workshop is the overall role Wages per worker higher by how much?


1
Expert's answer
2022-01-17T16:01:43-0500

Calculate mean

Workshop A


"mean_1=\mu_1"


"=\dfrac{1}{2+18+50+14+8+20}(2(2.50)+18(3.50)"

"+50(4.00)+14(3.00)+8(3.00)+20(2.00))=3.34"


Workshop B

"mean_2=\mu_2"

"=\dfrac{1}{12+7+10+6+2+1
}(12(3.00)+7(3.00)"

"+10(4.25)+6(3.50)+2(3.50)+1(5.00))=3.49"

Workshop B pays higher average rate of wages.


"3.49-3.34=0.15"